OpenStudy (anonymous):

Solve the system by substitution. 6x + 3y = -6 2x + y = -2

OpenStudy (anonymous):

6x + 3y = -6 => 3y = -6 - 6x => y = (-6-6x) / 3 Putting this value of y in equation 2 , we get : 2x + (-6-6x)/3 = -2 .. Solve it for x , and then for y ..